# Small maximal clusters are very unlikely in critical random graphs

### v1

@@ 1.4 Notation and structure of the paper
for some (new) constant $c=c(d)>0$. This concludes the proof of Theorem 1.1. What is left to do is giving proofs of Lemmas 2.6 and 2.7; this is done in the following section.
2.3.1 Proofs of Lemmas 2.6 and 2.7
In this section we establish the two auxiliary facts which were stated, without proof, while proving the main theorem for the $\mathbb{G}(n,d,p)$ model.
Proof of Lemma 2.6.
Recall that $\mathcal{G}_{2L}\coloneqq\{t_{2L}\leq BLT^{1/2}\}$. To bound from above the probability of the complementary event, we employ the same strategy used for the $\mathbb{G}(n,p)$ model. Let $S_{0}\coloneqq d$ and define recursively $S_{i}\coloneqq S_{i-1}+\hat{\eta}_{i}-1$, where the random variable $\hat{\eta}_{i}$ is defined (for $i\leq(dn)/2$) by
@@
Recalling (32) and (33) we immediately see that $\hat{\eta}_{i}=\eta_{i}$ when either $1\leq i\leq t_{1}$ or $i\in(t_{j-1}+1,t_{j}]$ for some $1<j\leq N$, whereas $\hat{\eta}_{t_{j-1}+1}=\eta_{t_{j-1}+1}-1$ for $1<j\leq N$. Indeed, if $1\leq i\leq t_{1}$, then $e_{i}$ is taken from the set of active half-edge, so that $\mathcal{S}(e_{i})\cap\mathcal{U}_{i-1}\setminus\{e_{i}\}=\mathcal{S}(e_{i})\cap\mathcal{U}_{i-1}=\emptyset$ (whence $|\mathcal{S}(e_{i})\cap\mathcal{U}_{i-1}\setminus\{e_{i}\}|=0$) and also $\mathcal{A}_{i-1}\cup\mathcal{S}(e_{i})=\mathcal{A}_{i-1},\mathcal{U}_{i-1}\setminus\mathcal{S}(e_{i})=\mathcal{U}_{i-1}$; thus $\hat{\eta}_{i}=\eta_{i}$. The other cases are treated similarly.
Then $(-\min_{j\in[t]}S_{j})+1$ denotes the number of disjoint clusters that have been fully explored by time $t\in[n]$. When $Y_{t}=0$, we have fully explored a cluster, and this occurs precisely when $S_{t}=\min_{j\in[t-1]}S_{j}-1$, i.e., the process $(S_{t})_{t\in[n]}$ reaches a new all-time minimum.
@@
$\mathbb{P}(\mathcal{G}_{2L}^{c})\leq\mathbb{P}\Big{(}\max_{j\leq\lfloor BLT^{1/2}\rfloor}(-S_{j})<2L\Big{)}.$ (40)
To continue, we separate $\hat{\eta}_{t}$ into its constituent by introducing
@@
$\mathbb{P}\Big{(}d+\sum_{i=1}^{j}X_{i,1}>-2L-\sum_{i=1}^{j}X_{i,2}\text{ }\forall j\leq\lfloor BLT^{1/2}\rfloor,\sum_{i=1}^{\lfloor BLT^{1/2}\rfloor}\left|\mathcal{S}(e_{i})\cap\mathcal{U}_{i-1}\setminus\{e_{i}\}\right|<2dL\Big{)}.$ (41)
@@
In order to control the (random) sum $\sum_{i=1}^{\lfloor BLT^{1/2}\rfloor}\mu^{\prime}_{i}$ we use Lemma $3.4$ in [13]
@@
Recalling that $L\sim An^{1/3},T\sim n^{2/3}/A$ and $p=(1+\lambda n^{-1/3})/(d-1)$ we obtain (for all large enough $n$)
$q(\lfloor BLT^{1/2}\rfloor)-5dL\geq L\Big{(}\frac{B^{2}LT(d-2)}{4nd(d-1)}-5d\Big{)}\geq L\Big{(}\frac{B^{2}(d-2)}{5d(d-1)}-5d\Big{)}.$
@@
$\lfloor BLT^{1/2}\rfloor\big{(}(d-1)^{-1}-p\big{)}\geq-|\lambda|BLT^{1/2}(d-1)^{-1}n^{-1/3}\geq-|\lambda|BL(d-1)^{-1}A^{-1/2}$
@@
$y\geq(d-1)^{-1}L\Big{(}\frac{B^{2}(d-2)}{5d(d-1)}-5d-|\lambda|BA^{-1/2}\Big{)}\geq c^{\prime}L$
@@
Since, by assumption, $A=O(n^{1/6})$, taking $m=L$ (which is indeed $O(n^{1/2})$) we see that the second and third terms on the right-hand side of the last expression are both at most $\exp(-\Theta(A^{3/2}))$, thus completing the proof of the lemma.
Proof of Lemma 2.7.
@@
$\mathbb{P}(t_{i}-t_{i-1}\geq T|\mathcal{F}_{i-1})\geq\frac{c_{0}}{T^{1/2}}$
for all large enough $n$, where we recall that $\mathcal{F}_{i-1}=\mathcal{F}_{t_{i-1}}$ is the $\sigma$-algebra consisting of all the information gathered by the exploration process until time $t_{i-1}$ (i.e. until the moment at which we start revealing the $(i-1)$-th cluster). Then, setting $Y^{i}_{t}\coloneqq Y_{t_{i}+t}$ and $\eta^{i}_{t}\coloneqq\eta_{t_{i}+t}$ to simplify notation, we can write
$\mathbb{P}(t_{i}-t_{i-1}\geq T|\mathcal{F}_{i-1})=\mathbb{P}(Y^{i}_{t}>0\text{}\forall t\leq T-1|\mathcal{F}_{i-1})\geq\mathbb{P}\Big{(}d_{i}+\sum_{k=1}^{t}\eta^{i}_{k}>0\text{ }\forall t\leq T|\mathcal{F}_{i-1}\Big{)},$ (50)
where $d_{i}\geq 1$ represents the number of unseen stubs attached to the node from which we start exploring (at time $t_{i-1}+1$) the $i$-th cluster. Recall that, if $Y^{i}_{k-1}\geq 1$, then
@@
where denotes the half-edge selected at time $t_{i-1}+k$ whereas $\mathcal{U}^{i}_{k-1},\mathcal{A}^{i}_{k-1}$ are the set of unseen stubs and set of active half-edges, respectively. We further denote by $\mathcal{V}^{(d)}_{k,i}$ the set of (fresh) vertices with $d$ unseen stubs at time $t_{i-1}+k$ and $F^{i}_{k}\coloneqq\{v(h^{i}_{k})\in\mathcal{V}^{(d)}_{k-1,i}\}$, the event that $v(h^{i}_{k})$ is fresh.
@@
Indeed, if at step $t_{i-1}+k$ we pick $h^{i}_{k}$ from the set of active vertices, then $\eta^{i}_{k}=\delta^{i}_{k}$; similarly, if $h^{i}_{k}$ is incident to a fresh vertex (whence in particular $h^{i}_{k}$ is unseen), then again $\eta^{i}_{k}=\delta^{i}_{k}$. In all other cases (i.e. when $h^{i}_{k}$ is unseen but the vertex to which is attached has less than $d$ unseen half-edges incident to it), then $\eta^{i}_{k}\geq\delta^{i}_{k}$.
The reason why the $\delta^{i}_{k}$ are a good approximation of the $\eta^{i}_{k}$ is because, during the first phase of the exploration process, most vertices are fresh (and hence for most steps $i\leq T$ we expect that $\eta^{i}_{k}=\delta^{i}_{k}$).
@@
$\mathbb{P}\Big{(}d_{i}+\sum_{k=1}^{t}\delta^{i}_{k}>0\text{ }\forall t\leq T|\mathcal{F}_{i-1}\Big{)}.$ (54)
To establish the required lower bound for this probability, we proceed exactly as in Section $4.2$ of [13]. Actually, in our case the proof is even simpler, because we are not conditioning on the event $\mathbb{S}_{n}$ that the multigraph $\mathbb{G}^{\prime}(n,d)$ underlying $\mathbb{G}^{\prime}(n,d,p)$ is simple (removing the conditioning on $\mathbb{S}_{n}$ was the content of Section $4.2.1$ in [13]). We will not give all the details here, but nevertheless, for the sake of clarity, we will carefully explain the required adaptations needed from [13] to the current setting.
The idea is to replace the $\delta^{i}_{k}$ with the simpler
@@
and then to show that, for any $\gamma\in(0,1/2)$, we can find a constant $c_{0}$ (which depends on $B,d$) such that, for all sufficiently large $A,n$, on $\mathcal{G}_{i-1}$ we can bound
$\mathbb{P}\Big{(}d_{i}+\sum_{i=1}^{t}\chi^{i}_{k}\geq t^{\gamma}\text{ }\forall t\leq T\mid\mathcal{F}_{i-1}\Big{)}\geq c^{\prime}_{0}T^{-1/2}$ (56)
for some constant $c^{\prime}_{0}=c^{\prime}_{0}(B,d)>0$. Subsequently we show that for $\gamma\in(0,1/2)$ (and for all large enough $A,n$), on $\mathcal{G}_{i-1}$ we have
@@
Let $\chi^{i}_{k}$ be as in (55). Suppose that (56) and (57) hold. Then, there exists a constant $c_{0}=c_{0}(B,d)>0$ such that, for all sufficiently large $A,n$, on $\mathcal{G}_{i-1}$ we have
$\mathbb{P}\Big{(}d_{i}+\sum_{k=1}^{t}\eta^{i}_{k}>0\text{ }\forall t\leq T|\mathcal{F}_{i-1}\Big{)}\geq c_{0}T^{-1/2},$
@@
We follow the argument given in Section $4.2.6$ in [13]. Note that
$\mathbb{P}\Big{(}d_{i}+\sum_{k=1}^{t}\delta^{i}_{k}>0\text{ }\forall t\leq T|\mathcal{F}_{i-1}\Big{)}=\mathbb{P}\Big{(}d_{i}+\sum_{k=1}^{t}\delta^{i}_{k}>0\text{ }\forall t\leq(t_{i}-t_{i-1})\wedge T|\mathcal{F}_{i-1}\Big{)}$ (58)
@@
$\mathbb{P}\Big{(}d_{i}+\sum_{k=1}^{t}\chi^{i}_{k}>t^{\gamma}\text{ }\forall t\leq(t_{i}-t_{i-1})\wedge T,\sum_{i=1}^{t}\mathbbm{1}_{\{h^{i}_{k}\in\mathcal{A}^{i}_{k-1}\}}<t^{\gamma}\text{ }\forall t\leq(t_{i}-t_{i-1})\wedge T\mid\mathcal{F}_{i-1}\Big{)},$
@@
There remains to establish (56) and (57). To this end, we continue with the following Lemma, whose proof follows the exact same steps as Proposition $4.11$ in [13].
@@
Define $\Delta^{i}_{k}\coloneqq\mathbbm{1}_{R^{i}_{k}}\mathbbm{1}_{\{U_{k}\leq 1-(4BLT^{1/2})/n\}}$ for $k\leq T$, where $(U_{k})_{k\in\mathbb{N}}$ is a sequence of iid $\text{Unif}([0,1])$ random variables, independent of everything else. Then
$\mathbb{P}\Big{(}\sum_{k=1}^{t}\chi^{i}_{k}\geq t^{\gamma}\text{ }\forall t\leq T\mid\mathcal{F}_{i-1}\Big{)}\geq\mathbb{P}\Big{(}\sum_{i=1}^{t}\Delta^{i}_{k}\geq t^{\gamma}\text{ }\forall t\leq T\Big{)}.$
The intuition here is that, since $|\mathcal{V}^{(d)}_{k,i}|\geq n-1-2(t_{i-1}+k)$ for each $k\geq 1$ (as at most two vertices can be removed from the set of fresh nodes at each step), the (conditional) probability (on the event $\mathcal{G}_{i-1}$) that $h^{i}_{k}$ is fresh when $k\leq T$ equals
@@
provided that $n$ is large enough. Thus (with $\mathcal{F}_{i-1}(k-1)\coloneqq\mathcal{F}_{t_{i-1}+k-1}$) for all large enough $n$ we obtain
@@
which then (recalling the definitions of $\chi^{i}_{k}$ and $\Delta^{i}_{k}$) shows that $\chi^{i}_{k}$ stochastically dominates $\Delta^{i}_{k}$ for each $k$.
We do not give the actual proof of Lemma 2.12 because, as we said earlier, it can be established in the exact same way as the proof of Proposition $4.11$ in [13]. (We remark that the $1-(4BLT^{1/2})/n$ in the definition of $\Delta^{i}_{k}$ given here translates into $1-T^{\prime}/n$ in Proposition $4.11$ of [13].)
@@
Let $\Delta^{i}_{k}$ be as in Lemma 2.12. Let $(D_{k})_{k\leq T}$ be a sequence of iid random variables with $\mathbb{P}(D_{k}=d-2)=1/(d-1)=1-\mathbb{P}(D_{k}=-1)$. There exists constants $A_{0}=A_{0}(\lambda,d)>0$ and $c=c(d)>0$ such that, for any $A\geq A_{0},\varepsilon\in(0,1)$ and for all large enough $n$, we have
@@
$\frac{\text{d}\hat{\mathbb{P}}}{\text{d}\mathbb{P}}\Big{|}_{\mathcal{F}_{t}}\coloneqq\frac{\exp\big{(}\nu\sum_{k=1}^{t}\Delta^{i}_{k}\big{)}}{\mathbb{E}\big{[}\exp\big{(}\nu\sum_{k=1}^{t}\Delta^{i}_{k}\big{)}\big{]}},\text{ }t\in\mathbb{N}_{0}$
@@
Thanks to (60), we only need to check that
@@
whence, recalling that $T\sim n^{2/3}/A$ and $L\sim An^{1/3}$, we see that
$T\Big{(}\frac{\lambda^{2}}{n^{2/3}}+\frac{\lambda BLT^{1/2}}{n^{4/3}}+\frac{(BLT^{1/2})^{2}}{n^{2}}\Big{)}\asymp\frac{\lambda^{2}}{A}+\frac{\lambda}{A^{1/2}}+1.$
Therefore $\mathbb{E}[e^{\nu\Delta^{i}_{1}}]^{T}\geq c$. Moreover,
@@
and so (since $A\ll n^{2/3}$) we see that $e^{-|\nu|T^{1/2}\varepsilon^{-1}}\geq c>0$ too, as required. ∎
@@
Combining Lemmas 2.12 and 2.13 together with (62) we conclude that indeed (56) holds. There remains to establish (57); this is the content of the next lemma, which is established following the steps carried out in Section $4.2.5$ of [13].
@@
$\mathbb{P}\Big{(}\exists t\leq(t_{i}-t_{i-1})\wedge b:h^{i}_{k}\in\mathcal{A}^{i}_{k-1}\mid\mathcal{F}_{i-1}\Big{)}\leq\sum_{k=1}^{b}\frac{d(k-1)}{dn-2(t_{i-1}+k-1)-1}\leq\frac{2b^{2}}{n}$
for all large enough $n$, where for the last inequality we have used that, on $\mathcal{G}_{i-1}$, we have $t_{i-1}+k\leq BLT^{1/2}+T\leq 2BLT^{1/2}$ and $BLT^{1/2}\asymp A^{1/2}n^{2/3}\ll n$ (as $A\ll n^{2/3}$).
@@
$\mathbb{P}\Big{(}\sum_{i=1}^{(t_{i}-t_{i-1})\wedge T}\mathbbm{1}_{\{h^{i}_{k}\in\mathcal{A}^{i}_{k-1}\}}\geq b^{\gamma}\mid\mathcal{F}_{i-1}\Big{)}\\ \leq\mathbb{P}\Big{(}\big{\{}\sum_{i=1}^{(t_{i}-t_{i-1})\wedge T}\mathbbm{1}_{\{h^{i}_{k}\in\mathcal{A}^{i}_{k-1}\}}\geq b^{\gamma}\big{\}}\cap\mathcal{B}\mid\mathcal{F}_{i-1}\Big{)}+\mathbb{P}(\mathcal{B}^{c}\mid\mathcal{F}_{i-1}).$ (64)
@@
($\mathbb{P}\Big{(}\big{\{}\sum_{i=1}^{(t_{i}-t_{i-1})\wedge T}\mathbbm{1}_{\{h^{i}_{k}\in\mathcal{A}^{i}_{k-1}\}}\geq b^{\gamma}\big{\}}\cap\mathcal{B}\mid\mathcal{F}_{i-1}\Big{)}\leq\exp\big{(}-b^{\gamma}+cT\ell n^{-1}\big{)}$

### v3

@@ 1.4 Some useful tools
for some (new) constant $c=c(d)>0$. This concludes the proof of Theorem 1.1. What is left to do is giving proofs of Lemmas 2.3 and 2.4; this is done in the following section.
2.3.1 Proofs of Propositions 2.3 and 2.4
In this section we establish the two auxiliary facts which were stated, without proof, while proving the main theorem for the $\mathbb{G}(n,d,p)$ random graph.
Proof of Proposition 2.3.
Recall that $\mathcal{G}_{2L-1}\coloneqq\{t_{2L-1}\leq BLT^{1/2}\}$. To bound from above the probability of the complementary event, we employ the same strategy used for the $\mathbb{G}(n,p)$ model. Let $S_{0}\coloneqq d$ and define recursively $S_{i}\coloneqq S_{i-1}+\hat{\eta}_{i}$, where the random variable $\hat{\eta}_{i}$ is defined (for $i\leq(dn)/2$) by
@@
Recalling (32) and (33) we immediately see that $\hat{\eta}_{i}=\eta_{i}$ when either $1\leq i\leq t_{1}$ or $i\in(t_{j-1}+1,t_{j}]$ for some $1<j\leq N$, whereas $\hat{\eta}_{t_{j-1}+1}=\eta_{t_{j-1}+1}-1$ for $1<j\leq N$. Indeed, if $1\leq i\leq t_{1}$, then $e_{i}$ is taken from the set of active half-edge, so that $\mathcal{S}(e_{i})\cap\mathcal{U}_{i-1}\setminus\{e_{i}\}=\mathcal{S}(e_{i})\cap\mathcal{U}_{i-1}=\emptyset$ (whence $|\mathcal{S}(e_{i})\cap\mathcal{U}_{i-1}\setminus\{e_{i}\}|=0$) and also $\mathcal{A}_{i-1}\cup\mathcal{S}(e_{i})=\mathcal{A}_{i-1},\mathcal{U}_{i-1}\setminus\mathcal{S}(e_{i})=\mathcal{U}_{i-1}$; thus $\hat{\eta}_{i}=\eta_{i}$. The other cases are treated similarly. Then $(-\min_{j\in[t]}S_{j})+1$ denotes the number of disjoint clusters that have been fully explored by time $t\in[n]$. When $Y_{t}=0$, we have fully explored a cluster, and this occurs precisely when $S_{t}=\min_{j\in[t-1]}S_{j}-1$, i.e., the process $(S_{t})_{t\in[n]}$ reaches a new all-time minimum.
@@
$\mathbb{P}(\mathcal{G}_{2L-1}^{c})\leq\mathbb{P}\Big{(}\max_{j\leq\lfloor BLT^{1/2}\rfloor}(-S_{j})<2L-1\Big{)}.$ (40)
To continue, we split $\hat{\eta}_{t}$ into its contributions by introducing
@@
$\mathbb{P}\Big{(}d+\sum_{i=1}^{j}X_{i,1}>-2L+1-\sum_{i=1}^{j}X_{i,2}\text{ }\forall j\leq\lfloor BLT^{1/2}\rfloor,\sum_{i=1}^{\lfloor BLT^{1/2}\rfloor}\left|\mathcal{S}(e_{i})\cap\mathcal{U}_{i-1}\setminus\{e_{i}\}\right|<2dL\Big{)}.$ (41)
@@
In order to control the (random) sum $\sum_{i=1}^{\lfloor BLT^{1/2}\rfloor}\mu^{\prime}_{i}$ we use Lemma $3.4$ in [13], which we recall here for the reader’s convenience. Note that the $m$ which appears in the statement comes from the definition of $\mu^{\prime}_{i}$.
@@
Recalling that $L\sim dAn^{1/3}/4(d-1),T\sim(d-1)n^{2/3}/A$ and $p=(1+\lambda n^{-1/3})/(d-1)$ we obtain (for all large enough $n$)
$q(\lfloor BLT^{1/2}\rfloor)-5dL\geq L\Big{(}\frac{B^{2}LT(d-2)}{4nd(d-1)}-5d\Big{)}\geq L\Big{(}\frac{B^{2}(d-2)}{18(d-1)}-5d\Big{)}.$
@@
$\lfloor BLT^{1/2}\rfloor\big{(}(d-1)^{-1}-p\big{)}\geq-|\lambda|BLT^{1/2}(d-1)^{-1}n^{-1/3}\geq-2|\lambda|BL(d-1)^{-1/2}A^{-1/2}$
@@
$y\geq(d-1)^{-1}L\Big{(}\frac{B^{2}(d-2)}{18(d-1)}-5d-2(d-1)^{1/2}|\lambda|BA^{-1/2}\Big{)}\geq c^{\prime}L$
@@
Since, by assumption, $A=O(n^{1/6})$, taking $m=L$ (which is indeed $O(n^{1/2})$ by our assumption on $A$) we see that the second and third terms on the right-hand side of the last inequality are both at most $\exp(-\Omega(A^{3/2}))$, thus completing the proof of the proposition.
Proof of Proposition 2.4.
@@
$\mathbb{P}(t_{i}-t_{i-1}\geq T|\mathcal{F}_{t_{i-1}})\geq\frac{c_{0}}{T^{1/2}}$
for all large enough $n$, where we recall that $\mathcal{F}_{t_{i-1}}$ is the $\sigma$-algebra consisting of all the information gathered by the exploration process until time $t_{i-1}$ (i.e. until the moment at which we finish revealing the $(i-1)$-th cluster). Then, setting $Y^{i}_{t}\coloneqq Y_{t_{i-1}+t}$ and $\eta^{i}_{t}\coloneqq\eta_{t_{i-1}+t}$ to simplify notation, we can write
$\mathbb{P}(t_{i}-t_{i-1}\geq T|\mathcal{F}_{t_{i-1}})=\mathbb{P}(Y^{i}_{t}>0\text{ }\forall t\leq T-1|\mathcal{F}_{t_{i-1}})\geq\mathbb{P}\Big{(}d_{i}+\sum_{k=1}^{t}\eta^{i}_{k}>0\text{ }\forall t\leq T|\mathcal{F}_{t_{i-1}}\Big{)},$ (50)
where $1\leq d_{i}\geq d$ represents the number of unseen stubs attached to the node from which we start exploring (at time $t_{i-1}+1$) the $i$-th cluster. Recall that, if $Y^{i}_{k-1}\geq 1$, then
@@
where we denote by $h^{i}_{k}$ the half-edge selected at time $t_{i-1}+k$, whereas $\mathcal{U}^{i}_{k-1},\mathcal{A}^{i}_{k-1}$ are the sets of unseen stubs and active half-edges at time $t_{i-1}+k-1$, respectively. We further denote by $\mathcal{V}^{(d)}_{k,i}$ the set of (fresh) vertices with $d$ unseen stubs at time $t_{i-1}+k$ and by $F^{i}_{k}\coloneqq\{v(h^{i}_{k})\in\mathcal{V}^{(d)}_{k-1,i}\}$ the event that $v(h^{i}_{k})$ is fresh.
@@
Indeed, if at step $t_{i-1}+k$ we pick $h^{i}_{k}$ from the set of active vertices, then $\eta^{i}_{k}=\delta^{i}_{k}$; similarly, if $h^{i}_{k}$ is incident to a fresh vertex (whence in particular $h^{i}_{k}$ is unseen), then again $\eta^{i}_{k}=\delta^{i}_{k}$. In all other cases (i.e. when $h^{i}_{k}$ is unseen but the vertex to which is incident has less than $d$ unseen half-edges incident to it), then $\eta^{i}_{k}\geq\delta^{i}_{k}$.
@@
$\mathbb{P}\Big{(}d_{i}+\sum_{k=1}^{t}\delta^{i}_{k}>0\text{ }\forall t\leq T|\mathcal{F}_{t_{i-1}}\Big{)}.$ (54)
To establish the required lower bound for this probability, we proceed exactly as in Section $4.2$ of [13]. Actually, in our case the proof is even simpler, because we are not conditioning on the event $\mathbb{S}_{n}$ that the multigraph $\mathbb{G}^{\prime}(n,d)$ underlying $\mathbb{G}^{\prime}(n,d,p)$ is simple (removing the conditioning on $\mathbb{S}_{n}$ was the content of Section $4.2.1$ in [13]). We do not give all the details here but nevertheless, for the sake of clarity, we carefully explain the required adaptations needed from [13] to the current setting.
First of all note that the probability in (54) is at least
$\mathbb{P}\Big{(}d_{i}+\sum_{k=1}^{t}\delta^{i}_{k}>t^{\gamma}\text{ }\forall t\leq T|\mathcal{F}_{t_{i-1}}\Big{)},$
where $20<\gamma<1/$ 2 (for $\gamma\geq 1/2$, the lower bound that we obtain is not of the required order). Then the idea is to replace the $\delta^{i}_{k}$ with the simpler random variables
@@
and then to show that, for any $\gamma\in(0,1/2)$, we can find a constant $c_{0}$ (which depends on $B,d$) such that, for all sufficiently large $A,n$, on the event $\mathcal{G}_{i-1}$ we can bound
$\mathbb{P}\Big{(}d_{i}+\sum_{i=1}^{t}\chi^{i}_{k}\geq t^{\gamma}\text{ }\forall t\leq T\mid\mathcal{F}_{t_{i-1}}\Big{)}\geq c^{\prime}_{0}T^{-1/2}$ (56)
for some constant $c^{\prime}_{0}=c^{\prime}_{0}(B,d)>0$. Subsequently we show that for $\gamma\in(0,1/2)$ (and for all large enough $A,n$), on the event $\mathcal{G}_{i-1}$ we have
@@
Let $\chi^{i}_{k}$ be as in (55). Suppose that (56) and (57) hold. Then, there exists a constant $c_{0}=c_{0}(B,d)>0$ such that, for all sufficiently large $A,n$, on the event $\mathcal{G}_{i-1}$ we have
$\mathbb{P}\Big{(}d_{i}+\sum_{k=1}^{t}\eta^{i}_{k}>0\text{ }\forall t\leq T|\mathcal{F}_{t_{i-1}}\Big{)}\geq c_{0}T^{-1/2},$
@@
We closely follow the argument given in Section $4.2.6$ in [13]. Note that
$\mathbb{P}\Big{(}d_{i}+\sum_{k=1}^{t}\delta^{i}_{k}>0\text{ }\forall t\leq T|\mathcal{F}_{t_{i-1}}\Big{)}=\mathbb{P}\Big{(}d_{i}+\sum_{k=1}^{t}\delta^{i}_{k}>0\text{ }\forall t\leq(t_{i}-t_{i-1})\wedge T|\mathcal{F}_{t_{i-1}}\Big{)}$ (58)
@@
$\mathbb{P}\Big{(}d_{i}+\sum_{k=1}^{t}\chi^{i}_{k}>t^{\gamma}\text{ }\forall t\leq(t_{i}-t_{i-1})\wedge T,\sum_{i=1}^{t}\mathbbm{1}_{\{h^{i}_{k}\in\mathcal{A}^{i}_{k-1}\}}<t^{\gamma}\text{ }\forall t\leq(t_{i}-t_{i-1})\wedge T\mid\mathcal{F}_{t_{i-1}}\Big{)},$
@@
There remains to establish (56) and (57). To this end, we continue with the following Lemma, whose proof follows the exact same steps as Proposition $4.11$ in [13] and is therefore omitted.
@@
Define $\Delta^{i}_{k}\coloneqq\mathbbm{1}_{R^{i}_{k}}\mathbbm{1}_{\{U_{k}\leq 1-(4BLT^{1/2})/n\}}(d-1)-1$ for $k\leq T$, where $(U_{k})_{k\in\mathbb{N}}$ is a sequence of iid $\text{Unif}([0,1])$ random variables, independent of everything else. Then
$\mathbb{P}\Big{(}\sum_{k=1}^{t}\chi^{i}_{k}\geq t^{\gamma}\text{ }\forall t\leq T\mid\mathcal{F}_{t_{i-1}}\Big{)}\geq\mathbb{P}\Big{(}\sum_{i=1}^{t}\Delta^{i}_{k}\geq t^{\gamma}\text{ }\forall t\leq T\Big{)}.$
The intuition here is that, since $|\mathcal{V}^{(d)}_{k,i}|\geq n-1-2(t_{i-1}+k)$ for each $k\geq 1$ (as at most two vertices can be removed from the set of fresh nodes at each step), the (conditional) probability that $h^{i}_{k}$ is fresh when $k\leq T$ is, on the event $\mathcal{G}_{i-1}$, equal to
@@
provided that $n$ is large enough. Thus, setting $\mathcal{F}_{i-1}(k-1)\coloneqq\mathcal{F}_{t_{i-1}+k-1}$, for all large enough $n$ we obtain
@@
which then (recalling the definitions of $\chi^{i}_{k}$ and $\Delta^{i}_{k}$) shows that $\chi^{i}_{k}$ stochastically dominates $\Delta^{i}_{k}$ for each $k$. (We remark that the $1-(4BLT^{1/2})/n$ in the definition of $\Delta^{i}_{k}$ given here translates into $1-T^{\prime}/n$ in Proposition $4.11$ of [13].)
@@
Let $\Delta^{i}_{k}$ be as in Lemma 2.8. Let $(D_{k})_{k\leq T}$ be a sequence of iid random variables with $\mathbb{P}(D_{k}=d-2)=1/(d-1)=1-\mathbb{P}(D_{k}=-1)$. There exist constants $A_{0}=A_{0}(\lambda,d)>0$ and $c=c(d)>0$ such that, for any $A\geq A_{0},\varepsilon\in(0,1)$ and for all large enough $n$, we have
@@
$\frac{\text{d}\hat{\mathbb{P}}}{\text{d}\mathbb{P}}\Big{|}_{\mathcal{F}_{t}}\coloneqq\frac{\exp\big{(}\nu\sum_{k=1}^{t}\Delta^{i}_{k}\big{)}}{\mathbb{E}\big{[}\exp\big{(}\nu\sum_{k=1}^{t}\Delta^{i}_{k}\big{)}\big{]}},\text{ }t\in\mathbb{N}_{0},$
@@
There remains to check that
@@
whence, recalling that $T\sim(d-1)n^{2/3}/A$ and $L=\Theta(An^{1/3})$, we see that
$T\Big{(}\frac{\lambda^{2}}{n^{2/3}}+\frac{\lambda BLT^{1/2}}{n^{4/3}}+\frac{(BLT^{1/2})^{2}}{n^{2}}\Big{)}=\Theta\big{(}\frac{\lambda^{2}}{A}+\frac{\lambda}{A^{1/2}}+1\big{)}.$
Therefore $\mathbb{E}[e^{\nu\Delta^{i}_{1}}]^{T}\geq c$ provided that $A\geq A_{0}$ for a large enough $A_{0}=A_{0}(\lambda,d)>0$. Moreover,
@@
and so (since $A\ll n^{2/3}$) we see that $e^{-|\nu|T^{1/2}\varepsilon^{-1}}\geq c>0$ too (provided that $A\geq A_{0}$ for a large enough $A_{0}=A_{0}(\lambda,d)>0$), as required. ∎
@@
Combining Lemmas 2.8 and 2.9 together with (62) we conclude that indeed (56) holds. There remains to establish (57); this is the content of the next lemma, which is established following the steps carried out in Section $4.2.5$ of [13].
@@
$\mathbb{P}\Big{(}\exists t\leq(t_{i}-t_{i-1})\wedge T:\sum_{i=1}^{t}\mathbbm{1}_{\{h^{i}_{k}\in\mathcal{A}^{i}_{k-1}\}}\geq t^{\gamma}\mid\mathcal{F}_{t_{i-1}}\Big{)}\ll T^{-1/2}.$ (63)
@@
$\mathbb{P}\Big{(}\exists t\leq(t_{i}-t_{i-1})\wedge b:h^{i}_{k}\in\mathcal{A}^{i}_{k-1}\mid\mathcal{F}_{t_{i-1}}\Big{)}\leq\sum_{k=1}^{b}\frac{d(k-1)}{dn-2(t_{i-1}+k-1)-1}\leq\frac{2b^{2}}{n}$
for all large enough $n$, where for the last inequality we have used that, on the event $\mathcal{G}_{i-1}$, we have $t_{i-1}+k\leq BLT^{1/2}+T\leq 2BLT^{1/2}$ and $BLT^{1/2}=\Theta(A^{1/2}n^{2/3})\ll n$ (as $A\ll n^{2/3}$).
@@
$\mathbb{P}\Big{(}\sum_{i=1}^{(t_{i}-t_{i-1})\wedge T}\mathbbm{1}_{\{h^{i}_{k}\in\mathcal{A}^{i}_{k-1}\}}\geq b^{\gamma}\mid\mathcal{F}_{t_{i-1}}\Big{)}\\ \leq\mathbb{P}\Big{(}\big{\{}\sum_{i=1}^{(t_{i}-t_{i-1})\wedge T}\mathbbm{1}_{\{h^{i}_{k}\in\mathcal{A}^{i}_{k-1}\}}\geq b^{\gamma}\big{\}}\cap\mathcal{B}\mid\mathcal{F}_{i-1}\Big{)}+\mathbb{P}(\mathcal{B}^{c}\mid\mathcal{F}_{t_{i-1}}).$ (64)
@@
($\mathbb{P}\Big{(}\big{\{}\sum_{i=1}^{(t_{i}-t_{i-1})\wedge T}\mathbbm{1}_{\{h^{i}_{k}\in\mathcal{A}^{i}_{k-1}\}}\geq b^{\gamma}\big{\}}\cap\mathcal{B}\mid\mathcal{F}_{t_{i-1}}\Big{)}\leq\exp\big{(}-b^{\gamma}+cT\ell n^{-1}\big{)}$